OpenStudy (anonymous):

enrollment at a particualr university is 11,600 there are seven times as many students who work full time as there are students who do not. How many students work full time and how many dont?

OpenStudy (anonymous):

Let's say that the number of students who don't work is "x". If that's true, then the number of students who do work can be represented by "7x". Since both the working and non-working number of students add up to 11600, we can set up the equation: 7x + x = 116000. Add like terms: 8x = 116000 And solve for x: x = 116000/8 = 1450 students who don't work. Plugging this x value back into "7x" we can find the number of students who do work: 7(1450)= 10150